OpenStudy (anonymous):

Find the monthly payment for the loan. (Round your answer to the nearest cent.) A $124,000 home bought with a 20% down payment and the balance financed for 30 years at 9.5%

OpenStudy (dangerousjesse):

I'm sorry about that :)

OpenStudy (anonymous):

Wow why would someone spew such things

OpenStudy (dangerousjesse):

We've been having problems with hackers and spammers lately, it'll go away eventually.

OpenStudy (anonymous):

that wasnt me :(

OpenStudy (anonymous):

Thanks you.

OpenStudy (dangerousjesse):

Ye okay

OpenStudy (anonymous):

but ill help you anyways

OpenStudy (dangerousjesse):

\(124,000\times .20=x\) \(124,000-x=y\) \(30\times 12=z\) \(y\div z=\textsf{the answer}\)

OpenStudy (anonymous):

Thank you. This is very helpful DangerousJesse